OpenStudy (anonymous):

Eliminate the "z" variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = -16 2x - y + 2z = -15 2x - 2y - z = 1

OpenStudy (anonymous):

x-4y=1 4x-5y=-13

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

a. 4x - 3y = -14 b. 4x + y = -14 c. 6x + 5y = 13 d. 6x - 5y = -13 Not an option :/ Only those are options

OpenStudy (anonymous):

hold on